OpenStudy (anonymous):

need the answer and fast will fan and medal he area of a triangle is 45.5 cm2 and the base is 10 cm. What is the height of the triangle?

rishavraj (rishavraj):

\[Area = \frac{ 1 }{ 2 } \times base \times height \]

OpenStudy (anonymous):

i know i dont know the answer and that does not help

OpenStudy (anonymous):

well you can plug the given information into the formula 45.5=(1/2)10h

rishavraj (rishavraj):

so \[h = \frac{ 45.5 \times 2 }{ 10 }\]

rishavraj (rishavraj):

@BrieiaKlarc done?????

OpenStudy (anonymous):

yea its 9.1

rishavraj (rishavraj):

yup ......